what is another solution to 144 besides 12 squared
Answer by Edwin McCravy(20054)


144 = 1×144
144 = 2×72
144 = 3×48
144 = 4×36
144 = 6×24
144 = 8×18
144 = 9×16
144 = 12×12 = 122
144 = 16×9
144 = 18×8
144 = 24×6
144 = 36×4
144 = 48×3
144 = 72×2
144 = 144×1
Edwin